A 150-g piece of iron at 3000C is submerged in 250 g of water at 100C to be cooled quickly.  

         Determine the final temperature of the iron and the water. [ 5 marks]

[ Cw  = 4.18x103 J/kg0C and Ciron= 4.5 x 102J/kg0C]


1
Expert's answer
2022-01-16T13:09:16-0500

From thermal equilibrium equation:


"Q_i+Q_w=0,\\
c_im_i(t_{f}-t_{ii})+c_wm_w(t_f-t_{iw})=0,\\\space\\
t_f=\dfrac{c_im_it_{ii}+c_wm_wt_{iw}}{c_im_i+c_wm_w}=27.6°\text{ C}."